Use the given information to write the equation of each quadratic equation. The intercepts of the parabola are (-1,0), (5,0) and (0,15).
My work,
-1 and 5 are the roots
while 15 is the y-int.
(x+1)(x-5)
x^2-4x-5
-3(x^2-4x-5) = -3x^2+12x+15

Answer by lwsshak3(11628)

Use the given information to write the equation of each quadratic equation. The intercepts of the parabola are (-1,0), (5,0) and (0,15).
My work,
-1 and 5 are the roots
while 15 is the y-int.
(x+1)(x-5)
x^2-4x-5
-3(x^2-4x-5) = -3x^2+12x+15
***
here is another way you can do it:
........y=ax^2+bx+c
(-1,0)..0=a-b+c
(5,0)...0=25a+5b+c
(0,15)..15=0+0+c
..
c=15
a-b=15
25a+5b+15=0
..
5a-5b+75=0
25a+5b+15=0
add
30a+90=0
30a=-90
a=-3
b=a+15=-3+15=12
..
a=-3, b=12, c=15
Equation: -3x^2+12x+15